the diameter of wire X is twice that of wire Y and tension in X is 4 times that of Y the ratio of velocity of waves in X and Y is

Grade:12th pass

1 Answers

Norman Manchu
6 years ago
Wave velocity though can be given by the ratio of w by k it can also be expressed as the root of ratio of trtension and mass Here's how!
\large v=^{\sqrt{\frac{T}{m}}}=^{\sqrt{\frac{T}{ \rho V}}}=^{\sqrt{\frac{T}{ \rho \pi r^{2}l}}}
Thus v1|v2=√T1x(r2)^2|√T2x(r1)^2=1:1